OpenStudy (littleheron):

Calculate the slope of the points (-2, -1) and (3, 4). a. 1 b. 2 c. 3 d. 4

OpenStudy (♪chibiterasu):

Slope Formula \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (♪chibiterasu):

Plug in and solve. :)

OpenStudy (♪chibiterasu):

(x, y) <-- for reference

OpenStudy (littleheron):

so its B?

OpenStudy (♪chibiterasu):

Not quite, try calculating again.

OpenStudy (♪chibiterasu):

\[\frac{ 4 - (-1) }{ 3 - (-2) }\]

OpenStudy (♪chibiterasu):

- ( - just turns into + so \[\frac{ 4 + 1 }{ 3 + 2 }\]

OpenStudy (littleheron):

that would be 5?

OpenStudy (♪chibiterasu):

Both of them would be 5, so now divide

OpenStudy (♪chibiterasu):

What happens when you divide a number by itself?

OpenStudy (littleheron):

you get 1

OpenStudy (♪chibiterasu):

=)

OpenStudy (♪chibiterasu):

So just remember that formula and calculating slope with only two points should be easy.

OpenStudy (littleheron):

so its a?

OpenStudy (♪chibiterasu):

yeap

OpenStudy (littleheron):

cool thanks!